3)Find the area of an equilateral triangle with side 14.

Answer by Boreal(15235)

Area is (1/2) bh
base is 14
half the base is 7
the altitude is a 30-60-90 triangle where the short leg is 7, so the long leg is 7 sqrt (3)
Area is (1/2)7*7*sqrt (3)= (49/2) sqrt (3)=42.44 square units.
A= side^2 *(sqrt (3)/4)
=49*sqrt (3)/2, which is the same result